OpenStudy (anonymous):

Hi someone could help me with the following differential equation excata, thank (cosy+ycosx)dx+(senx-xseny)dy=0

OpenStudy (amistre64):

sen?

OpenStudy (anonymous):

sin

OpenStudy (anonymous):

(cosy+ycosx)dx+(sinx-xsiny)dy=0

OpenStudy (anonymous):

excuse me

OpenStudy (amistre64):

well, exacts have a nice property .... inegrate the left wrt.x what do we get?

OpenStudy (fozia):

u can apply ecaxt differntial eqation

OpenStudy (amistre64):

\[\int \cos y+y~\cos x~dx\]

OpenStudy (anonymous):

yes

OpenStudy (fozia):

let M=(cosy+ycosx) and N=senx-xseny...now take partialy derivat of M w.r.to y and take partialy dervtof N w r to x

OpenStudy (amistre64):

"yes" is not the integration

OpenStudy (anonymous):

Ok

OpenStudy (fozia):

this gives you My=-siny+cosx and Nx=cosx-siny that both the derivativs are equal.now use integrl of M with r to x+int of N (only y terms)=0 gives you result

OpenStudy (anonymous):

ok thank you very much, thank you for your help